A boat travels 45 mi upstream(against the current) in 5 h.The boat travels the same distance downstream in 3 h.What is the rate of the boat in still water?What is the rate of the current?
Answer by Paul(988)

Let the speed of the boat be y
Let the speed of the current be x
5(x-y)=45 ---> solve for y
5x-5y=45
-5y=45-5x
y=9+x
Subsitution

3(x+y)=45
3x+3y=45 (allow subsitution)
3x+3(9+x)=45
3x+27+3x=45
6x=18
x=3

y=9+3
y=12

Hence, the speed of the current is 3mph and the speed of the boat is 12mph.
Paul.